OpenStudy (anonymous):

A projectile is thrown upward so that its distance above the ground after t seconds is h = - 16t2 + 480t. After how many seconds does it reach its maximum height?

OpenStudy (anonymous):

Take the derivative and find where it is equal to 0.

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

\[\frac{ dh }{ dt }=-32t+480\] \[0=-32t+480\] \[-32t=-480 \rightarrow 32t=480\] \[t=\frac{ 480 }{ 32 }\]

OpenStudy (anonymous):

so the answer would be 15s

OpenStudy (anonymous):

yes